1. Mrs. Salcedo's car of mass 2.4 x 10³ kg travelling on the highway when the teacher receives an emergency call. She increased the speed of the car to 33 m/s.The increase in speed results in 3.1 x10^5 J of work done on the car. Determine the initial speed of the car in km/h.



1
Expert's answer
2021-11-21T17:25:09-0500

"mv^2/2-mv_0^2/2=W\to v_0=\sqrt{\frac{2}{m}(\frac{mv^2}{2}-W)}="


"=\sqrt{\frac{2}{2400}\cdot(\frac{2400\cdot 33^2}{2}-310000)}=11.4\ (m/s)\approx41\ (km/h)"